AO SDT on 08 November 2011 UT

Observer(s):C. Arcidiacono, F. Manucci, J. Argamedo, E. Pinna, A.Skemer,JC. Guerra (LBTO)

Telescope Operator: D. Huerta (LBTO)

SW Support: T. Sargent (Tucson)

Telescope Support: M. Mirkeff (LBTO)

Instrument Support: J. Morris (LBTO)

Summary

Dome not open. Lost the night.
Not observing due to bad weather and ice on the top. 
Humidity:98%
Temperature outside:-8
Retro-reflector ON.

Test of close loop with the retro-reflector. 
Mount of the ND filter holder.

Details

Check of the observinfg scripts for the next nights. Check the observing scripts sending presets to telescope and to the AO system. Mount of the ND filter holder in the Pisces buffle. The plate to fit the filter holder was modified to give some movements freeedom. It was seeing that placing the plate all the way to the left and up the "vignetting" is less. A work for tomorrow afternoon is to check this and to improve the center positon.

-- JuanGuerra - 08 Nov 2011
Topic revision: r4 - 13 Nov 2011, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback